Twilight:

I need help: The two figures are similar. Write a proportion to find the missing measure. Then find the value of x.

MxxnLight:

@jhonyy9

Twilight:

jhonyy9:

so bc given that these figures are similar you can use the proportionality of corresponding sides and write in this way

jhonyy9:

\[\frac{ 66 }{ 11 } = \frac{ x }{ 9 }\] cross multiplie and calcule the x value

jhonyy9:

@MxxnLight

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 \[\frac{ 66 }{ 11 } = \frac{ x }{ 9 }\] cross multiplie and calcule the x value \(\color{#0cbb34}{\text{End of Quote}}\) x = 66*9/11 = ? @Laylalyssa